You have just received an offer in the mail from Friendly Loans. The company is offering to loan you $5,250 with low monthly payments of $95 per month. If the interest rate on the loan is an APR of 14.6 percent compounded monthly, how long will it take for you to pay off the loan?
Answer #1

use financial calculator as

PV=-5250 is the loan amount

FV=0 as loan is repaid

PMT=95 is monthly repayment of loan

1/Y=14.6/12=1.2167 is monthly rate

Click CPT

Click N=92.27 months or 7.69 years

the above is the answer